y''-2y'-3y=2e-x-10sinx


1
Expert's answer
2021-08-09T13:09:37-0400

Let us solve the differential equation "y''-2y'-3y=2e^{-x}-10\sin x." The characteristic equation "k^2-2k-3=0" is equivalent to "(k+1)(k-3)=0," and hence has the roots "k_1=-1,k_2=3."

The general solution of the differential equation is "y=C_1e^{-x}+C_2e^{3x}+y_p," where the particular solution is "y_p=axe^{-x}+b\cos x+c\sin x." Then "y'_p=ae^{-x}-axe^{-x}-b\sin x+c\cos x, y''_p=-ae^{-x}-ae^{-x}+axe^{-x}-b\cos x-c\sin x."

It follows that

"-2ae^{-x}+axe^{-x}-b\cos x-c\sin x-2(ae^{-x}-axe^{-x}-b\sin x+c\cos x)-3(axe^{-x}+b\cos x+c\sin x)=2e^{-x}-10\sin x."

Then

"-4ae^{-x}+(-4b-2c)\cos x+(2b-4c)\sin x=2e^{-x}-10\sin x."

We conclude that "-4a=2,-4b-2c=0, 2b-4c=-10." It follows that "a=-\frac{1}{2},c=-2b, 10b=-10." Therefore, "a=-\frac{1}{2}, b=-1,c=2."

We conclude that the general solutiion of the differential equation is

"y=C_1e^{-x}+C_2e^{3x}-\frac{1}{2}xe^{-x}-\cos x+2\sin x."